OpenStudy (anonymous):

What is the radius of a spherical balloon with a volume of 36 m3?

Directrix (directrix):

We need the volume of a sphere formula.

Directrix (directrix):

V = (4/3) * pi* r^3

Directrix (directrix):

36 = (4/3) * pi * r^3 27 = pi* r^3

Directrix (directrix):

@princessemilyhehe Solve the above equation for r.

OpenStudy (anonymous):

radius is 2.048352189765887 rounded is 2.05